Prepare (using a spreadsheet package) an entire duration amortization schedule as follows: Initial amount: Shs 4,000,000/=; rate of interest: 13.5% per annum; period: 10 years payable monthly in arrears. An additional amount of Shs 400,000/= is repaid with the 30th installment, and the monthly installment recomputed to fully pay the loan balance in the initially agreed period. The rate of interest is adjusted to 15.5% per annum with effect from the end of the fourth year. The monthly repayment is recomputed to fully pay the loan balance in the initially agreed period. The borrower enhances his monthly repayments by Shs 12,000/= per month from the end of month 74. The borrower is advanced an additional Shs 400,000/= at the end of month 80 and the monthly repayment is recomputed to fully pay the loan balance in 9 months less than had initially been agreed
Answer #1
Pv Initial amount          4,000,000
Rate Monthly interest =13.5/12= 1.125%
Nper Number of months =10*12 120
PMT Monthly payment          60,909.72 (Using PMT function of excel with Rate =1.125%, Nper=120,Pv=-4000000)
As per Mortgage Schedule for Month 1-30,
PV1 Loan Balance at end of 30 months    3,036,016.37
Rate Monthly interest =13.5/12= 1.125%
Nper1 Number of months left(120-30) 90
PMT1 Monthly payment from 31st month to 120th month          53,818.98 (Using PMT function of excel with Rate =1.125%, Nper=90,Pv=-3036016.37)
As per Mortgage Schedule for Month 31-48,
Pv2 Loan Balance at end of 48 months    2,646,102.53
Rate2 Monthly interest rate =15.5/12= 1.292% 1.2916667%
Nper2 Number of months left(120-48) 72
PMT2 Monthly payment from 48 month to 120th month          56,672.87 (Using PMT function of excel with Rate =1.292%, Nper=72,Pv=-2646102.53)
Pv3 Loan Balance at end of 80 months    1,274,363.86
Rate Monthly interest rate =15.5/12= 1.292%
Nper Number of months left(120-80-9)) 31
PMT3 Monthly payment from 48 month to 111th month          50,148.01 (Using PMT function of excel with Rate =1.292%, Nper=31,Pv=-1274363.86)

Clipboard Font Alignment Number Styles * v fac =PMT(F4,F5,-F3) G H I J K L M N Pv Rate Initial amount Monthly interest =13.5/MORTGAGE SCHEDULE MONTH1-30 Month Beginning Monthly Balance Payment 1 4,000,000 60,909.72 2 3,984,090.28 60,909.72 3 3,968,00MORTGAGE SCHEDULE FROM MONTH 31-48 C=A*1.125% D=B-C E=A-D Month from Beginning Month 31 33 А Beginning Monthly Balance PaymenMORTGAGE SCHEDULE FROM MONTH 49-73 C=A*1.292% D=B-C E=A-D Month from Beginning Month 49 Beginning Balance 1 2,646,102.53 2 2,25 2,017,915.45 56,672.87 2 6,064.74 3 0,608.13 1,987,307.32 MORTGAGE SCHEDULE FROM MONTH 74-80 26 27 28 29 30 31 32 1,987,30MORTGAGE SCHEDULE Month 81-111 C=A1.2922 D=B-C E-A-D Month from Beginning Month Beginning Monthly Balance Payment 1 1,274,36